1.     Calculate the freezing and boiling points of a solution prepared by dissolving   

     15.5 g of Al (NO3)3 in 200.0 g of water. (Molar mass of Al (NO3)3 is 212.996 g/mol). 


1
Expert's answer
2021-04-29T07:23:30-0400

The reaction takes place as:


"Al(NO_3)_3\Rightarrow Al^{3+}+3{NO_3}^-"


Now here the mole of Aluminium Nitrate dissociates in 4 moles ,hence the value of (i=4)


Formula used "\Rightarrow" delta "T_f=i\times K_F \times m"

where "K_f=" 1.86 degre celcius kg "mol^{-1}"

m"=" molality


Now :

Moles of aluminium nitrate="\dfrac{weight}{molar weight}"

= "\dfrac{15.5}{212.996}=0.0728" moles


Hence :

Molality(m)="\dfrac{moles}{weight of solvent(gm)}\times1000"

m="\dfrac{0.0728}{200}\times1000=0.364\dfrac{mol}{gm}"

Now: For Freezing point:

delta "T_f=i\times K_f \times m"

"=4\times 1.86 \times0.364=2.70" degre celcius


Now we know that

delta "T_f={T^0}_f-T_f"

Putting "{T}^0_f=0"

We get the value of "T_f=-2.70" degree celcius and this is our answer.


Now for the boiling point

delta"T_b=i\times K_b \times m"

"=" "4\times 0.512 \times0.364=0.74" degre celcius

Now:

delta "T_b=T_b-{T}^0_b"

"T_b=0.714+100=100.714" degree celcius is our answer